A gas mixture contains equal no. of molecules of N2 and SF6, some of it is passed thro gaseous effusion apparatus. Calcutlate how many molecules of N2 are present in product gas for every 100 molecules of SF6.

Grade:11

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi Vishrant,

M1=28

M2=140

M2/M1=5

r1/r2=√5=2.236

x1/x2=r1/r2=2.236

x1=2.236*100≈223